### v1

@@ 1 Introduction
where $\theta^{\prime}=-\theta$ and $\gamma^{\prime}=\mu-\gamma$ are tailored scaling parameters. This work [38] not only establishes the exponential decay rate for the continuous level but also develops several accelerated primal-dual methods based on proper implicit, semi-implicit and explicit numerical schemes. Motivated by this, we introduce the following accelerated multiobjective primal-dual gradient (AMPD) flow for solving ˜LCMOP:
@@ 2.1 Notation
The collection of all continuous differentiable functions from $0,\infty)$ to ${\mathbb{R}}^{d}$ is denoted by the set $C^{1}([0,\infty);\,{\mathbb{R}}^{d})$, and $AC([0,\infty);\,{\mathbb{R}}^{d})$ consists of all absolutely continuous functions from $0,\infty)$ to ${\mathbb{R}}^{d}$. Let $\mathcal{F}_{L}^{1}$ be the set of all $C^{1}$ convex functions on $\mathbb{R}^{d}$ with $L$-Lipschitz continuous gradients. For any $f\in\mathcal{F}_{L}^{1}$ we have $\left\lVert{\nabla f(x)-\nabla f(y)}\right\rVert\leq L\left\lVert{x-y}\right\rVert$ and
@@
All $C^{1}$ functions on ${\mathbb{R}}^{d}$ that are $\mu$-strongly convex with some $\mu\geq 0$ constitute another important function class $\mathcal{S}_{\mu}^{1}$:
@@
For later use, define $\mathcal{S}^{1,1}_{\mu,L}:=\mathcal{S}_{\mu}^{1}\cap\mathcal{F}_{L}^{1}.$ The following result is trivial (cf. [44, Lemma 2.1]) but provides a useful property for the subsequent convergence rate analysis.
@@ 2.4 New merit function
The nonnegativity has been verified by the above discussions, and by [10, Lemmas 1.26 and 1.29], $\Pi$ is lower semicontinuous since $f_{j}$ is closed (i.e. lower semicontinuous) for all $1\leq j\leq m$.
@@
    Under ˜2, for any $x\in\mathcal{L}_{F}(\bm{\alpha})$, we have
@@
Notice that $\mathcal{L}_{F}(\bm{\alpha})=\cap_{j=1}^{m}\mathcal{L}_{f_{j}}(\bm{\alpha}_{j})$. According to ˜1, the level set $\mathcal{L}_{f_{j}}(\bm{\alpha}_{j})$ is bounded for all $1\leq j\leq m$ with $R_{j}(\bm{\alpha}_{j})<+\infty$. Then for all $z\in P_{w}\cap\mathcal{L}_{F}(\bm{\alpha})$, we have $\|z\|\leq\min_{1\leq j\leq m}R_{j}(\bm{\alpha}_{j})<+\infty$. This verifies the first claim (i).
@@
By ˜2, for all $z\in\mathcal{L}_{F}(\bm{\alpha})\cap\Omega$, there exists $z^{*}\in P_{w}$ such that $F(z^{*})\leq F(z)$. This implies the following identity
@@
Assume that $f_{j}\in\mathcal{F}^{1}_{L_{j}}$ for $1\leq j\leq m$ and $K\subset\,{\mathbb{R}}^{n}$ is bounded with ${\bf diam}\left({K}\right)$$\infty$. Then for any $x\in K$ we have
@@ 3.1 Continuous model
The nonnegative constant $\mu\geq 0$ is the minimal convexity parameter among all objectives and will be clarified later in ˜3. Clearly, the case $\beta=0$ is allowed, which reduces to ˜1.
@@ 3.3 Lyapunov analysis
for all $t>0$. In our subsequent analysis, we need the following assumption.
Assumption 3.
Assume $f_{j}\in\mathcal{S}_{\mu_{j},L_{j}}^{1,1}$ for $1\leq j\leq m$ with $0\leq\mu_{j}\leq L_{j}<+\infty$. For simplicity,we also denote $\mu:=\min_{1\leq j\leq m}\mu_{j}$ and $L:=\max_{1\leq j\leq m}L_{j}$.
@@
Since $q\in C(x)={\bf conv}\left\{{\nabla f_{1}(x),\cdots,\nabla f_{m}(x)}\right\}$, assume that $q=\sum_{j=1}^{m}\lambda_{j}\nabla f_{j}(x)$ with some $\lambda\in\Delta_{m}$. Then it follows from ˜3 and 3 that
@@ 4 Rate of Convergence
where $\bm{\alpha}(\bar{x})=\max\{\bm{\alpha}_{*},F(x_{0})+C_{1}(R(\alpha_{0}(\bar{x})))\}$. Above, the constant $\bm{\alpha}_{*}$ has been declared in ˜2, and $R(\cdot),\,D(\cdot)$ and $C_{0}(\cdot,\cdot)$ are defined respectively in ˜1, 8 and 25.
@@
Similarly with ˜32, we can prove that $f_{j(t)}(\bar{x})-f_{j(t)}(x_{0})\leq C_{1}(\left\lVert{\bar{x}}\right\rVert)$, which leads to $f_{j(t)}(x(t))\leq\alpha_{0}(\bar{x})$ and thus $x(t)\in\mathcal{L}_{f_{j(t)}}(\alpha_{0}(\bar{x}))$. By ˜1 we conclude that $\left\lVert{x(t)}\right\rVert\leq R(\alpha_{0}(\bar{x}))$ for all $t>0$. This completes the proof. ∎
@@
According to ˜2, $\mathcal{L}_{F}(\bm{\alpha}_{*})\cap\Omega$ is nonempty and so is $\mathcal{L}_{F}(\bm{\alpha}(\bar{x}))\cap\Omega$ since $\bm{\alpha}(\bar{x})=\max\{\bm{\alpha}_{*},F(x_{0})+C_{1}(R(\alpha_{0}(\bar{x})))\}$. Noticing that $x(t)\in\ \mathcal{L}_{F}({\bm{\alpha}}(\bar{x}))$, by using Lemma˜2.4 we obtain the upper bound estimate
@@ 5.1 Numerical scheme
Algorithm 1 AMPD-QP for solving ˜LCMOP with $f_{j}\in\mathcal{S}_{\mu_{j},L_{j}}^{1,1}$
@@ 5.2 Discrete Lyapunov analysis
Let $\lambda_{k}=(\lambda_{k,1},\cdots,\lambda_{k,m})^{\top}\in\Delta_{m}$ be such that the projection admits the presentation ${\bf proj}_{C(y_{k})}\big(w_{k+1}-A^{\top}\widehat{\xi}_{k}\big)=\sum_{j=1}^{m}\lambda_{k,j}\nabla f_{j}(y_{k})$. Then we obtain from Lemma˜2.1 that
@@ 7 Conclusion
In this work, we develop a novel continuous-time primal-dual framework for solving ˜LCMOP. Based on a new merit function, we introduce the concept of an $\epsilon$-approximation solution to the weakly Pareto optimality. We then propose an accelerated multiobjective primal-dual flow and establish the exponential decay via the Lyapunov analysis. In addition, we consider an implicit-explicit discretization scheme and prove that both the feasibility violation and the objective gap have the same rates $\mathcal{O}(1/k)$ and $\mathcal{O}(1/k^{2})$ respectively for the convex case and the strongly convex case.

### v2

@@ 1 Introduction
where $\theta^{\prime}=-\theta$ and $\gamma^{\prime}=\mu-\gamma$ are tailored scaling parameters and $\mu\geq 0$ is the convexity parameter of the objective $f$. This work [38] not only establishes the exponential decay rate for the continuous level but also develops several accelerated primal-dual methods based on proper implicit, semi-implicit and explicit numerical schemes. Motivated by this, we introduce the following accelerated multiobjective primal-dual (AMPD) flow for solving ˜LCMOP:
@@ 2.1 Notation
The collection of all continuous differentiable functions from $0,\infty)$ to ${\mathbb{R}}^{d}$ is denoted by the set $C^{1}([0,\infty);\,{\mathbb{R}}^{d})$, and $AC([0,\infty);\,{\mathbb{R}}^{d})$ consists of all absolutely continuous functions from $0,\infty)$ to ${\mathbb{R}}^{d}$. Let $\mathcal{F}_{L}^{1}(\,{\mathbb{R}}^{d})$ be the set of all $C^{1}$ convex functions on $\mathbb{R}^{d}$ with $L$-Lipschitz continuous gradients. For any $f\in\mathcal{F}_{L}^{1}(\,{\mathbb{R}}^{d})$ we have $\left\lVert{\nabla f(x)-\nabla f(y)}\right\rVert\leq L\left\lVert{x-y}\right\rVert$ and
@@
All $C^{1}$ functions on ${\mathbb{R}}^{d}$ that are $\mu$-strongly convex with some $\mu\geq 0$ constitute another important function class $\mathcal{S}_{\mu}^{1}(\,{\mathbb{R}}^{d})$:
@@
For later use, define $\mathcal{S}^{1,1}_{\mu,L}(\,{\mathbb{R}}^{d}):=\mathcal{S}_{\mu}^{1}(\,{\mathbb{R}}^{d})\cap\mathcal{F}_{L}^{1}(\,{\mathbb{R}}^{d})$. The result given below is trivial (cf. [44, Lemma 2.1]) but provides a useful property for the subsequent convergence rate analysis.
@@
Throughout, we impose the following assumption on the objectives of ˜LCMOP. Note that in our setting, the minimal strong convexity constant is not necessarily positive. In other words, we focus on not only the convex case $\mu=0$ but also the strongly convex case $\mu>0$, in a unified way.
Assumption 1.
Assume $f_{j}\in\mathcal{S}_{\mu_{j},L_{j}}^{1,1}(\,{\mathbb{R}}^{n})$ for $1\leq j\leq m$ with $0\leq\mu_{j}\leq L_{j}<+\infty$. For simplicity, we also denote $\mu:=\min_{1\leq j\leq m}\mu_{j}$ and $L:=\max_{1\leq j\leq m}L_{j}$.
@@ 2.4 New merit function
The nonnegativity has been verified by the above discussions, and by [10, Lemmas 1.26 and 1.29], $\Pi$ is lower semicontinuous since $f_{j}$ is continuous differentiable and as well as closed (i.e. lower semicontinuous) for all $1\leq j\leq m$.
@@
    Under ˜3, for any $x\in\mathcal{L}_{F}(\bm{\alpha})$, we have
@@
Notice that $\mathcal{L}_{F}(\bm{\alpha})=\cap_{j=1}^{m}\mathcal{L}_{f_{j}}(\bm{\alpha}_{j})$. According to ˜2, the level set $\mathcal{L}_{f_{j}}(\bm{\alpha}_{j})$ is bounded for all $1\leq j\leq m$ with $R_{j}(\bm{\alpha}_{j})<+\infty$. Then for all $z\in P_{w}\cap\mathcal{L}_{F}(\bm{\alpha})$, we have $\|z\|\leq\min_{1\leq j\leq m}R_{j}(\bm{\alpha}_{j})<+\infty$. This verifies the first claim (i).
@@
By ˜3, for all $z\in\mathcal{L}_{F}(\bm{\alpha})\cap\Omega$, there exists $z^{*}\in P_{w}$ such that $F(z^{*})\leq F(z)$. This implies the following identity
@@
Let $K\subset\,{\mathbb{R}}^{n}$ be such that ${\bf diam}\left({K}\right)$$\infty$. Then for any $x\in K$ we have
@@ 3.1 Continuous model
The nonnegative constant $\mu\geq 0$ is the minimal convexity parameter of all objectives and has been clarified in ˜1. Clearly, the case $\beta=0$ is allowed, which reduces to ˜1.
@@ 3.3 Lyapunov analysis
for all $t>0$. In our subsequent analysis, we need the following lemma, which tells us how to calculate the derivative of the Lyapunov function with respect to the time variable.
@@
Since $q\in C(x)={\bf conv}\left\{{\nabla f_{1}(x),\cdots,\nabla f_{m}(x)}\right\}$, assume that $q=\sum_{j=1}^{m}\lambda_{j}\nabla f_{j}(x)$ with some $\lambda\in\Delta_{m}$. Then it follows from ˜3 and 1 that
@@ 4 Rate of Convergence
where $\bm{\alpha}(\bar{x})=\max\{\bm{\alpha}_{*},F(x_{0})+C_{1}(R(\alpha_{0}(\bar{x})))\}$. Above, the constant $\bm{\alpha}_{*}$ has been declared in ˜3, and $R(\cdot),\,D(\cdot)$ and $C_{0}(\cdot,\cdot)$ are defined respectively in ˜2, 8 and 25.
@@
Similarly with ˜32, we can prove that $f_{j(t)}(\bar{x})-f_{j(t)}(x_{0})\leq C_{1}(\left\lVert{\bar{x}}\right\rVert)$, which leads to $f_{j(t)}(x(t))\leq\alpha_{0}(\bar{x})$ and thus $x(t)\in\mathcal{L}_{f_{j(t)}}(\alpha_{0}(\bar{x}))$. By ˜2 we conclude that $\left\lVert{x(t)}\right\rVert\leq R(\alpha_{0}(\bar{x}))$ for all $t>0$. This completes the proof. ∎
@@
According to ˜3, $\mathcal{L}_{F}(\bm{\alpha}_{*})\cap\Omega$ is nonempty and so is $\mathcal{L}_{F}(\bm{\alpha}(\bar{x}))\cap\Omega$ since $\bm{\alpha}(\bar{x})=\max\{\bm{\alpha}_{*},F(x_{0})+C_{1}(R(\alpha_{0}(\bar{x})))\}$. Noticing that $x(t)\in\ \mathcal{L}_{F}({\bm{\alpha}}(\bar{x}))$, by using Lemma˜2.4 we obtain the upper bound estimate
@@ 5.1 Numerical scheme
Algorithm 1 AMPD-QP for solving ˜LCMOP with $f_{j}\in\mathcal{S}_{\mu_{j},L_{j}}^{1,1}(\,{\mathbb{R}}^{n})$
@@ 5.2 Discrete Lyapunov analysis
Let $\lambda_{k}=(\lambda_{k,1},\cdots,\lambda_{k,m})^{\top}\in\Delta_{m}$ be such that the projection admits the presentation ${\bf proj}_{C(y_{k})}\big(w_{k+1}-A^{\top}\widehat{\xi}_{k}\big)=\sum_{j=1}^{m}\lambda_{k,j}\nabla f_{j}(y_{k})$. Then we obtain from Lemmas˜2.1 and 1 that
@@ 7 Conclusion
In this work, we develop a novel continuous-time primal-dual framework for ˜LCMOP. Based on a new merit function, we introduce the concept of an $\epsilon$-approximation solution to the weakly Pareto optimality. We then propose an accelerated multiobjective primal-dual flow and establish the exponential decay via the Lyapunov analysis. In addition, we consider an implicit-explicit discretization scheme and prove that both the feasibility violation and the objective gap have the same rates $\mathcal{O}(1/k)$ and $\mathcal{O}(1/k^{2})$ respectively for the convex case and the strongly convex case.